OpenStudy (anonymous):

A boy can row a boat at a constant rate of 5 mi/hr in still water. He rows upstream for 15 minutes and then rows downstream returning to his starting point in another 10 minutes. Upstream net speed=5 -x mi/hr. downstresam net speed=5+x mi/hr. Find the rate of the current mi/hr and the total distance traveled

OpenStudy (anonymous):

d/(5-x) = 15 d = 15(5-x) d/(5+x) = 10 d = 10(5+x) you know how to do the rest.